A closed cylindrical tin is of height h cm and radius r cm, its total surface area is A cm2 and its volume is r cm3. Find an expression for A in terms of r . Taking , find an expression of v in terms of r , hence determine the value of r which make v maximum.
1
Expert's answer
2020-08-05T15:20:02-0400

Radius of cylinder = r cm

Height of cylinder = h cm

Volume of cylinder = πr²h cm³

But given that volume= r cm³

So πr²h = r

=> h = r/πr²

=> h = 1/πr

=> πrh = 1

Total surface area = 2πrh + 2πr²

So A = 2πrh + 2πr² = 2+ 2πr²

=> A = 2+ 2πr²

This is the expression for A in term of r

When A is constant,

A = 2πrh + 2πr²

=> 2πrh = A - 2πr²

So volume v = πr²h ="\frac{ r(A-2πr²)}{2}"

=> v = "\frac{Ar}{2}" - πr³

"So \frac {dv}{dr}" = "\frac {A}{2}" - 3πr²

For maximum value of v, "\frac {dv}{dr} =0"

=> "\frac {A}{2}" - 3πr² = 0

=> r² = "\frac {A}{6π}"

=> r = "\sqrt{\frac {A}{6π}}"

"\frac {d²v}{dr²}" = -6πr

Obviously "\frac {d²v}{dr²}" is negative when r = "\sqrt{\frac {A}{6π}}"

So v is maximum when r = "\sqrt{\frac {A}{6π}}"